What is the formula for solving this? "If you have a class consisting of 10 girls and 20 boys, in how many different ways can you form a committee from the class members such that there will be 2 girls and 2 boys on the committee?"
Answer by Theo(13342)

formula is:

c(10,2) * c(20,2)

that's (10! / (2!8!)) * (20!/(2!18!))

that becomes:

(10*9/2*1) * (20*19/2*1) which equals 45*190 = 8550

to see this in action, assume this was 3 girls and 4 boys and you want a committee of 2 girls and 2 boys.

number of possible combinations for girls is c(3,2) = 3!/2!1! = 3

assume the girls are abc.

then you get possible combinations of:

ab
ac
bc

anything else would be a duplicate. example: ba is the same combination of ab because order is not considered. if order was considered, it would be a permutation.

for the boys, the possible combinations would be c(4,2) = 4!/2!2! = 6

if the boys are defg, thenpossible combinations are:

de
df
dg
ef
eg
fg

now, for each combination of girls, you get 6 combinations of boys, which means a total of:

c(3,2)*c(4,2) = 3 * 6 = 18

these combinations would be:

ab + de
ab + df
ab + dg
ab + ef
ab + eg
ab + fg

ac + de
ac + df
ac + dg
ac + ef
ac + eg
ac + fg

bc + de
bc + df
bc + dg
bc + ef
bc + eg
bc + fg

same principle applies for the bigger figures, so your answer should be as stated above.